# From Pixels to Torques with Linear Feedback

### v2

@@ 5.1 Simulation Setup
Figure 4: Heat-map visualizations of each normalized row of the pixel-based Luenberger observer’s gain matrix, $L$. A cartpole visualization is also overlayed for reference. Each row of $L$ corresponds to the correction an image observation contributes to a respective state variable. Interestingly, visual features can be distinguished for each respective state variable: the cart for the cart position with the addition of the pole tip for the pole angle. The velocities also have similar features.
We evaluate the closed-loop performance of our pixel-based linear output-feedback policy in a simulated cartpole environment. We specify two models: a nominal model, which is simplified and contains up to $5\%$ parametric model error, and a true model, which is used exclusively for simulating the system in a “real-world” environment to evaluate the algorithm’s performance. We additionally introduce Gaussian process noise applied to the control inputs of the true model. The nominal model is used to design the teacher LQG policy, which uses encoder measurements of the cartpole’s configuration as observations for its state estimator. The encoder quantization is simulated as observation noise in the true model. To simulate a camera sensor for the pixels-to-torques student policy, we use Makie [9] to render greyscale images of the simulated cartpole. To discern these image-based observations from the true, “real-world” system, we visualize the true system in Meshcat [10].
(a) Cartpole-swing-up trajectory visualization. The goal configuration is shown in red.
@@
Figure 6: Overview of the cartpole hardware used for real-world demonstrations of pixels-to-torques control performed by linear output-feedback policies. The cartpole hardware is actuated by a brushless motor while an opposing global-shutter, Arducam OV2311 camera captures images of the cartpole. A Jetson Nano acts as the primary computer tasked with processing the images and executing the pixel-based student policy (or corresponding teacher) before sending the control inputs to the cartpole motor. This operation is performed at 60 Hz to match the camera’s maximum frame rate. Arrows indicate positive directions for the cart position, pole angle, and corresponding velocities.
@@
(b) Time histories of the cartpole configurations and corresponding control inputs. The teacher’s state estimates and anticipated control inputs are also shown as a reference.
Figure 7: A hardware demonstration of a pixel-based linear output-feedback policy stabilizing a real-world cartpole based on image-feedback from an Arducam OV2311 camera. Lightweight processing is performed on the camera images in the form of Gaussian blur and Otsu thresholding before being sent to the policy’s pixel-based Luenberger observer. Surprisingly, the pixel-based linear output-feedback policy is able to successfully stabilize the cartpole even under the presence of perturbations and visual occlusions.
@@ 5.2 Simulation Results
As shown in Fig. 4, we also investigate the observer gain of the pixel-based linear feedback policy. Specifically, we visualize each row of the policy’s observer gain, $L$, as a normalized heat map to discern the pixels that contribute most to the respective state variable’s estimation. Interestingly, doing so yields identifiable, best-corresponding visual features: Pixels that correspond to the cart contribute most to the cart-position estimation while the pole-tip pixels additionally contribute to the pole-angle estimation. In addition, the sign of the gain correctly corresponds to the physical location of the respective pixel. The respective velocities also have similar visual patterns.
We additionally showcase the Koopman-based extension of the pixel-based linear output-feedback policy for tracking a nonlinear system as shown in Fig. 5. Specifically, we demonstrate the Koopman-based policy’s ability to track a reference swing-up trajectory on the cartpole. The Koopman lifting consists of the cartpole states, 4th-order Fourier bases, and 6th-order Chebyshev polynomials. The reference trajectory was designed using iterative LQR before being tracked by a subsequent time-varying-LQG teacher policy for data collection. The Koopman-based policy is trained on 20 teacher demonstrations. As shown in Fig. 5(b), the policy is able to properly track the swing-up trajectory despite model mismatch and process noise.
@@ 5.3 Hardware Setup
We develop a physical cartpole system, shown in Fig. 6, to evaluate the efficacy of the pixel-based linear output-feedback policy on real-world hardware. The cartpole itself consists of a single brushless motor, a cart that rides along an aluminum track, and a pole laser cut from white acrylic. A black background is added to aid in image segmentation, and reflective components are also covered with a black cloth during runtime. Opposing the cartpole is a Jetson Nano and Arducam OV2311 camera mounted on a tripod. The camera has a global shutter to avoid rolling shutter distortion. The Jetson Nano acts as the main computer to perform online tasks, which consists of basic image processing (Gaussian blurring and Otsu thresholding) before executing the pixel-based linear output-feedback policy for state estimation and control. The system then sends control inputs to the motor at 60 Hz to match the frame rate of the camera. Image capture and processing is performed with OpenCV.
@@ 5.4 Hardware Results
As shown in Fig. 7, the pixels-to-torques policy is able to successful recover from perturbations and stabilize the real-world cartpole. Surprisingly, the policy is also robust to major occlusions, as shown in Fig. 7(a). While the occlusions do cause decreased state-estimation performance, as shown in Fig. 7(b), the policy’s state-feedback controller is still able to overcome the disturbance to stabilize the cartpole.

### v3

@@ 5.1 Simulation Setup
Figure 4: Heat-map visualizations of each normalized row of the pixel-based Luenberger observer’s gain matrix, $L$. A cartpole visualization is also overlayed for reference. Each row of $L$ corresponds to the correction an image observation contributes to a respective state variable. Interestingly, visual features can be distinguished for each state variable: the cart for the cart position with the addition of the pole tip for the pole angle. The velocities also have similar features.
We evaluate the closed-loop performance of our pixel-based linear output-feedback policy in a simulated cartpole environment. We specify two models: a nominal model, which is simplified and contains up to $5\%$ parametric model error, and a true model, which is used exclusively for simulating the system in a “real-world” environment to evaluate the algorithm’s performance. We additionally introduce Gaussian noise applied to the control inputs of the true model. The nominal model is used to design the teacher LQG policy, which uses encoder measurements of the cartpole’s configuration as observations for its state estimator. Observation noise is simulated as encoder quantization in the true model. To simulate a camera sensor for the pixels-to-torques student policy, we use Makie [9] to render greyscale images of the simulated cartpole at a resolution of $125\times 160$ pixels. To discern these image-based observations from the true, “real-world” system, we visualize the true system in Meshcat [10]. The pixels-to-torques student policy is learned on a computer equipped with a $64$-core AMD Threadripper CPU and $64$GB of RAM.
(a) Cartpole swing-up trajectory visualization. The goal configuration is shown in red.
@@
Figure 6: Overview of the cartpole hardware used for real-world demonstrations of pixels-to-torques control performed by linear output-feedback policies. The cartpole hardware is actuated by a brushless motor while an opposing global-shutter Arducam OV2311 camera captures images of the cartpole. A Jetson Nano acts as the primary computer tasked with processing the images and executing the pixel-based student policy (or corresponding teacher) before sending the control inputs to the cartpole motor. This operation is performed at 60 Hz to match the camera’s maximum frame rate. Arrows indicate positive directions for the cart position, pole angle, and corresponding velocities.
@@
(b) Time histories of the cartpole configurations and corresponding control inputs. The teacher’s estimates and anticipated control inputs are also shown as a reference.
Figure 7: A hardware demonstration of a pixel-based linear output-feedback policy stabilizing a real-world cartpole based on image-feedback from an Arducam OV2311 camera. Lightweight processing is performed on the camera images in the form of Gaussian blur and Otsu thresholding before being sent to the policy’s pixel-based Luenberger observer. Surprisingly, the pixel-based linear output-feedback policy is able to successfully stabilize the cartpole even under the presence of large perturbations and visual occlusions.
@@ 5.2 Simulation Results
As shown in Fig. 4, we also investigate the observer gain of the pixel-based linear feedback policy. Specifically, we visualize each row of the policy’s observer gain, $L$, as a normalized heat map to discern the pixels that contribute most to the respective state variable’s estimation. Interestingly, doing so yields identifiable visual features: Pixels that correspond to the cart contribute most to the cart-position estimate while the pole-tip pixels additionally contribute to the pole-angle estimate. In addition, the sign of the gain correctly corresponds to the physical location of the respective pixel.
We additionally showcase the Koopman-based extension of the pixel-based linear output-feedback policy for tracking a nonlinear system as shown in Fig. 5. Specifically, we demonstrate the Koopman-based policy’s ability to track a reference swing-up trajectory on the cartpole. The 76-dimensional Koopman lifting consists of the cartpole states, 4th-order Fourier features, and 6th-order Chebyshev polynomial features. The reference trajectory was designed using iterative LQR before being tracked by a time-varying-LQG teacher policy for data collection. The Koopman-based policy is trained on 20 demonstrations, each consisting of 75 samples over 2.5 seconds. As shown in Fig. 5(b), the policy is able to properly track the swing-up trajectory despite model mismatch and process noise.
@@ 5.3 Hardware Setup
We develop a physical cartpole system, shown in Fig. 6, to evaluate the efficacy of the pixel-based linear output-feedback policy on real-world hardware. The cartpole itself consists of a single brushless motor, a cart that rides along an aluminum track, and a pole laser cut from white acrylic. A black background is added to aid image contrast, and reflective components are also covered with a black cloth during runtime. Opposing the cartpole is a Jetson Nano and Arducam OV2311 camera mounted on a tripod. The camera has a global shutter to avoid rolling-shutter distortion. The Jetson Nano acts as the main computer to perform online tasks, which consists of basic image processing (Gaussian blurring and Otsu thresholding) before executing the pixel-based linear output-feedback policy for state estimation and control. The system then sends control inputs to the motor at 60 Hz to match the frame rate of the camera. Image capture and processing is performed with OpenCV at an image resolution of $115\times 140$ pixels. The pixels-to-torques student policy is trained with only 20 stabilizing demonstrations on the hardware, each with 150 samples over 2.5 seconds. The learning process is performed on a separate computer equipped with a 64-core AMD Threadripper CPU and 64GB of RAM, resulting in a solve time of about 4 minutes and 25 seconds.
@@ 5.4 Hardware Results
As shown in Fig. 7, the pixels-to-torques policy is able to successful recover from perturbations and stabilize the real-world cartpole. Surprisingly, the policy is also robust to major occlusions, as shown in Fig. 7(a). While the occlusions do cause decreased state-estimation performance, as shown in Fig. 7(b), the policy’s state-feedback controller is still able to overcome the disturbance to stabilize the cartpole. We attribute this to the sparse representation of $L$ as depicted in Fig. 4, which prevents a majority of the occluded pixels from affecting the policy’s internal state estimate.
@@ 6 Conclusions
Our linear pixels-to-torques approach has several limitations: First, the teacher-student methodology assumes that a successful teacher policy can be designed a priori with an appropriate internal state representation. Second, the method assumes un-obstructed observations during data collection. Third, by mapping raw pixel values directly to state estimates, our approach may also be sensitive to the camera’s calibration and pose. Finally, the nonlinear extension of our method also suffers from the limitations of Koopman-based approaches more broadly, which include the difficulty of crafting a good set of features or “observables” for lifting.